OpenStudy (anonymous):

(x-3)(3x-1)=(x+2)(x-1) Answer: 1/2, 5 How do i get there??

OpenStudy (anonymous):

multiply out first

OpenStudy (anonymous):

I'm trying to simplify.

OpenStudy (anonymous):

\[3x^2-10x+3=x^2+x-2\]

OpenStudy (anonymous):

then arrange as \[2x^2-11x+5=0\] and then solve

OpenStudy (anonymous):

\[(2x-1)(x-5)=0\] etc

OpenStudy (anonymous):

Thanks! ^_^